# ToDo (#1): #2

### v3

@@ 1.1.2 New Particles: Experimentally and Theoretically Motivated
Figure 1.4: Left: Example $X\rightarrow\nu\bar{\nu}$ signatures for 0.5% photon mixing with 8Li-produced $\bar{\nu}_{e}$ background (gray). Right: $X$ mass and $X$-to-photon production ratio 90% CL sensitivity. From Ref. [alonso_neutrino_2022].
@@
Another example of IsoDAR’s capability to look for new particles is the “$X$,” which is a low mass mediator that would couple to the photons produced in the target [alonso_neutrino_2022]. The $X$ would be a light new particle that decays to $\bar{\nu}\nu$ and would leave a distinct signature above background in the IBD spectrum. Fig.1.4 (left) displays resulting signatures for various masses (see color bar) assuming 0.5% $X$-to-photon production. The IBD background from $\bar{\nu}_{e}$ stemming from 8Li decays (gray) constitutes the primary background and has a distinct shape. The sensitivity is depicted in Fig.1.4 (right), where the blue region is excluded in the case of the SM. No existing limits are shown because this search is first-of-its-kind. A 17 MeV $X$ particle is consistent with the Atomki anomaly.
1.1.3 New Symmetries: Strong CP Violation and the QCD Axion
@@
Figure 1.5: IsoDAR sensitivity for axions (a) with mass $m_{a}$ coupling ($g$) to electrons (lines) [waites_axionlike_2023]. Gold dots–axion allowed; Teal–excluded regions; Gray–parameters with astrophysical implications.
@@
The neutron beam/bottle experiment lifetime discrepancy [Nico:2004ie, Yue:2013qrc, Serebrov:2017bzo, UCNt:2021pcg] motivates neutron–dark sector couplings. Through $n\rightarrow n^{\prime}\rightarrow n$ transitions, we may see neutrons disappearing from the target and reappear in the LSC. The underground environment provides a low background and the LSC’s size further adds to the sensitivity [Hostert:2022ntu]. Alternatively, if the $n^{\prime}$ are Majorana, then $n\rightarrow n^{\prime}\rightarrow\bar{n}$ may occur, depositing $\sim$2 GeV of energy in the LSC upon annihilation. This is a $B$ violating effect with implications for the early universe [Berezhiani:2020vbe].
The photons and electrons in the target may also couple to axions and IsoDAR can probe uncharted territory with astrophysical and cosmological implications [waites_axionlike_2023]. In Fig. 1.5 we show QCD axions coupling to electrons.
1.1.4 New Interactions: Tests via Precision Electroweak Physics
IsoDAR will record about 7000 elastic scatter events over its 4-year accumulated up-time. This constitutes the world’s largest sample at low momentum transfer. LEP and LHC data predict rate and energy dependency very precisely within the standard model [ParticleDataGroup:2020ssz]. With the normalization determined by the $\approx 1.7$M IBD events in the LSC, we will be able to measure $\sin^{2}\theta_{W}$ with a precision of 1.9% (see Fig.1.6 left).
As a neutrino-nucleon experiment, the observed deviation of NuTeV may stem from unaccounted-for nuclear physics effects; however, numerous possibilities have been investigated, but none fully explain the deviation. Alternatively, the deviation might signal new physics in the neutrino sector. If so, this BSM signature will also manifest in the IsoDAR ES measurement. Fig.1.6, right, presents two expectations for IsoDAR, where the red assumes background reduction from Cherenkov ring identification in LSC, while the green indicates the expectation without this additional information. Statistics limit the outcome, and further running can proceed with IsoDAR and DUNE as it comes online if a NuTeV-like deviation is observed.
@@ 1.1.4 New Interactions: Tests via Precision Electroweak Physics
Figure 1.6: Left: Weak Mixing Angle expectation for IsoDAR compared to present results and the future DUNE expectation. Right: expectation for IsoDAR in NSI parameter space. From Ref. [alonso_neutrino_2022].
@@
The accurately predicted energy dependence of the ES sample enables a precise phenomenological search for non-standard neutrino interactions (NSIs). This powerful search for lepton-lepton couplings complements the lepton-quark coupling searches by the active, global coherent neutrino scattering program. The SM right- and left-handed couplings, $g_{R}$ and $g_{L}$, are altered by NSI corrections for electron flavor antineutrino interactions with electrons (hence ”ee”) to $\tilde{g}_{R}=g_{R}+\varepsilon_{ee}^{eR}$ and $\tilde{g}_{L}=g_{L}+\varepsilon_{ee}^{eL}$. The SM expectation for IsoDAR@Yemilab in the parameter space of these deviations is seen in Fig. 1.6 (right), where red assumes background reduction using Cherenkov light, and green assumes only scintillation light reconstruction. This can be compared to the world result [guzzo] in black, which is dominated by $\nu_{e}$–electron ES scattering from beam dump experiments and, hence, is rotated with respect to IsoDAR, which has a pure $\bar{\nu}_{e}$ flux.
1.2 Requirements for the Proton Driver
@@
As shown in Fig. 1.1, the Proton Driver consists of a 60 MeV/amu cyclotron accelerating and extracting 5 electrical milliamperes of $\mathrm{H}_{2}^{+}$, a transport line that strips the $\mathrm{H}_{2}^{+}$ ions close to the extraction point, analyzes the stripped beam to monitor the health of the stripping foil, and transports the 10 milliamperes of protons to the target. The target consists of three nested hemispherical beryllium shells, the outermost being 20 cm in diameter, cooled with heavy water flowing around the shell structure. The proton beam is shaped, with quadrupoles and wobbler magnets to produce a distribution that allows for optimal heat transfer to the cooling water. Neutrons produced in the target system are moderated and flood into a roughly-spherical sleeve, about 1 meter in diameter, containing a mixture of beryllium chips and highly-enriched 7Li. The beta-decay of the 8Li resulting from the capture of neutrons produces the neutrinos of interest in the experiment.
@@
The principal requirements for the accelerator system are to reliably produce a 10 mA beam of protons in the most efficient manner possible with the lowest possible loss of particles in the acceleration process. Particularly above the Coulomb threshold. The experience of the cyclotron system at PSI [Stammbach] is that keeping the total power of beam loss inside the cyclotron vault to less than 200 W still allows access for component maintenance. Considering that the total beam power is 600 kW, one sees that this is a very stringent requirement.
As mentioned above, Volume I of this PDR addresses the physics and engineering aspects of the 60 MeV/amu cyclotron. The final chapter addresses the complex issue of transport and assembly of large and heavy cyclotron components in the underground environment at Yemilab.
A forthcoming Volume II will address the details of the beam transport system and the target, along with shielding requirements for the environment of the target, including the prevention of fast neutrons from reaching the fiducial volume of the liquid scintillator detector.
@@ 1.3 Overview of the Accelerator Layout
Figure 1.7: Overview of the components of the HCHC-60 Cyclotron System. Only one RF cavity is shown to not clutter the image.
To guide the reader through this PDR and, indeed, the path of an ion from genesis in the ion source to the full 60 MeV/amu and extraction, we provide an overview in Fig. 1.7. The main components of the system, which will be described in much detail in the following chapters, are:
  • The ion source, providing 10 mA of $\mathrm{H}_{2}^{+}$ at an emittance $<0.1\pi$-mm-mrad.
  • The Low Energy Beam Transport (LEBT), shaping and guiding the beam to the RFQ and providing diagnostics, and chopping for commissioning and personnel- and machine protection.
@@
Chapter 2 Design of the Front End
@@ 2.1 The Ion Source
The $\mathrm{H}_{2}^{+}$ beam to be injected into the cyclotron and accelerated to 60 MeV/amu will be produced by a filament-driven multicusp ion source. We have built a first prototype, the MIST-1 (Multicusp Ion Source at MIT - version 1), the technical details and first measured results of which are reported in Refs. [axani_high_2015, winklehner_first_2018, winklehner_high-current_2021-1, winklehner_new_2022, weigel_epics_2023]. Here we summarize these results and present ongoing upgrade work (MIST-2). The upgrade to MIST-2 improves on the MIST-1 by implementing several lessons learned during the systematic measurements and runtime experiences described below. A publication on the MIST-2 design and performance is forthcoming.
@@ 2.1.1 Operating Principle
In a multicusp ion source [ehlers:multicusp1], a plasma is confined by a set of permanent magnets that are arranged around the plasma chamber with alternating polarity, thereby creating “cusps” in the magnetic field. On average, the electrons and ions produced in the center of the chamber see an increasing magnetic field as they travel outward, which eventually turns them back. One notable exception are the cusps, which are the natural loss points of this configuration.
@@
Neutral gas is fed into the source to facilitate the generation of a high-density plasma. The type of gas determines the ion species produced in the source. Here we use H2 gas to form a plasma – and beam – comprising protons, $\mathrm{H}_{2}^{+}$, and H+${}_{3}^{+}$.
Table 2.1: A summary of the parameters for the MIST-1 ion source.
@@ 2.1.2 Mechanical Design
Figure 2.1: Cross section of MIST-1 (left) with extraction system and LEBT. Also shown is an example set of ion trajectories simulated with IBSimu and WARP. From Ref. [waites_high_2022].
For visual reference as we discuss the details of the ion source, 3D CAD renderings of MIST-1 (with subsequent LEBT) and MIST-2 are provided in Figs. 2.1 and 2.2, respectively. The main ion source parameters are also summarized in Table 2.1.
The MIST-type ion sources comprise a plasma chamber surrounded by a “water-jacket,” a back plate, and a front plate. The water jacket lets us pump de-ionized (DI) water around the plasma chamber for cooling. The radial confinement magnets are mounted inside the water-jacket and are cooled directly. At 6.5 cm, the plasma chamber is short compared to other ion sources. This is a direct consequence of the short mean free path of $\mathrm{H}_{2}^{+}$ in the plasma/neutral gas of the ion source discussed above.
@@
The front plate holds the front confinement magnets as well as the plasma aperture, a 75%W-25%Cu (aka “Elkonite”) plate that is embedded in the front plate. We have several plasma apertures with different-size openings ranging from 4 mm to 8 mm to test beam quality and total extractable current for various aperture diameters.
As seen in Figs. 2.1 and 2.2, following the plasma aperture is the “extraction system” which shapes and guides the beam as it is initially formed. The MIST-1 extraction system comprises five electrodes which shape and accelerate the ion beam. These include a plasma electrode (+15 kV), a puller electrode (-3 to -20 kV), and three electrostatic lenses – Lens 1 (+10 kV to +15 kV), lens 2 (-3 kV to -15 kV), and lens 3 (+10 kV to +15 kV).
Figure 2.2: MIST-2 CAD rendering calling out the most important features.
2.1.2.1 MIST-2 vs. MIST-1
@@
  • Replace all stainless steel components with OFHC copper to improve heat transfer
  • Increase the cross-section areas of all water cooling channels
@@
These changes will improve long-term running stability, reduce maintenance downtime, and allow for a better understanding of the plasma and beam conditions. Our high-fidelity simulations using IBSimu [kalvasIBSIMUThreedimensionalSimulation2010a] and WARP [vayNovelMethodsParticleInCell2012] suggest that the beam dynamics downstream of the six-way cross will not change between the MIST-1 and MIST-2 configurations. All particle simulations and designs still hold for LEBT, RFQ, and Cyclotron.

### v4

@@ 2.3.1 New Particles
Figure 5: Left: Example $X\rightarrow\nu\bar{\nu}$ signatures for 0.5% photon mixing with 8Li-produced $\bar{\nu}_{e}$ background (gray). Right: $X$ mass and $X$-to-photon production ratio 90% CL sensitivity. From Ref. [10].
Example 2: Production of a new particle coupling to photons—
An entirely new design enables searches for previously unknown particles. For an IsoDAR example, consider a new low mass mediator, which we will call “XX”, that mixes photons produced in the target and decays to $\bar{\nu}\nu$, where the $\bar{\nu}$ is observable through the IBD interaction [10].
@@
This analysis used the contained monoenergetic photon flux. The monoenergetic nature of the flux, hence the total energy of the $X$, leads to distinct peaks in the energy deposited by the $\bar{\nu}_{e}$ interactions. Fig. 5 (left) displays resulting signatures for various masses (see color bar) assuming 0.5% $X$-to-photon production. The IBD background from $\bar{\nu}_{e}$ stemming from 8Li decays (gray) constitutes the primary background and has a distinct shape. The sensitivity is depicted in Fig. 5 (right), where the blue region is excluded in the case of the SM. No existing limits are shown because this search is a first-of-its-kind.
2.3.2 New Symmetries
@@
Figure 6: IsoDAR sensitivity for axions (a) with mass $m_{a}$ coupling ($g$) to electrons (lines) [21]. Gold dots–axion allowed; Teal–excluded regions; Gray–parameters with astrophysical implications.
Example 1: A Dark Sector—
Dark sector models were initially motivated by the existence of dark matter, which cannot be accommodated within the SM. The model postulates a “dark sector” of particles and forces that interact with the SM through “portals.” Although motivated by the dark matter puzzle, dark sector models have expanded to address other theoretical and experimental puzzles. IsoDAR can search for evidence of the dark sector through mixing of the SM neutron, $n$, to a dark-sector neutron, $n^{\prime}$. Along with general connections to dark matter, by way of a dark sector, this effect is motivated by the discrepancy in neutron lifetime measurement between neutron beam and neutron bottle experiments [13, 14, 15, 16].
@@
This analysis utilizes the contained neutron flux. Through $n\rightarrow n^{\prime}\rightarrow n$ transitions, neutrons can disappear from the target and reappear in the LSC. Along with the essential aspects of the source design discussed above, the underground environment that provides a low background and the LSC’s size, allowing tight fiducial cuts add to the sensitivity [17]. Alternatively, if the $n^{\prime}$ are Majorana, then $n\rightarrow n^{\prime}\rightarrow\bar{n}$ may occur ($\bar{n}\ldotsanti$-neutron), depositing $\sim$2 GeV of energy in the LSC upon annihilation. This process is a $B$ violating effect with implications for the early universe [22]. This contribution to dark-sector studies is unique within a wide range of proposed dark-sector experiments, as discussed in Ref. [23].
Example 2: The QCD Axion—
The QCD axion arises in theories that introduce a new symmetry to explain why CP violation is suppressed in QCD [24, 25]. The introduction of the symmetry-breaking mechanism to make the new model parameter sufficiently small to agree with experimental data leads to the introduction of a new particle, the axion, $a$. There has been a long history of searches for the QCD axion at accelerators [25], and only a small open window of parameter space remains. IsoDAR can probe this territory [21].
This search employs the contained monoenergetic photon flux, generating axions by their coupling to photons. Axions can also couple to electrons, and the signature in the LSC is from interactions and decays involving both electrons and photons. In Fig. 6 we show QCD axions coupling to electrons as an example of the capability.
2.3.3 New Interactions
Example 1: The Weak Mixing Angle —
Despite many open questions within the SM, the theory is remarkably internally self-consistent within the electroweak sector [26]. As a result, LEP and LHC data can be used to predict the rate and energy dependence of neutrino scattering precisely. Deviations might indicate new physics. The NuTeV experiment observed a deviation of the weak mixing angle ($\sin^{2}\theta_{W}$) parameter from the SM prediction [27]. If this is new physics, it must be associated with neutrino properties, as polarized electron experiments have not observed the effect. However, as a neutrino-nucleon experiment, the NuTeV anomaly may stem from unaccounted-for nuclear physics effects; however, numerous possibilities have been investigated, but none fully explain the deviation and so the deviation remains unexplained [26].
IsoDAR offers a new approach to this question through measurement using neutrino-electron elastic scattering—a purely leptonic process that evades the concerns related to nuclear physics in NuTeV. IsoDAR will record $\sim$7000 ES events over its 4-year accumulated up-time. This constitutes the world’s largest sample, allowing the most precise measurement of $\sin^{2}\theta_{W}$ at low momentum transfer, $Q$. Fig.7, right, presents two expectations for IsoDAR, where the red assumes background reduction from Cherenkov ring identification in LSC, while the green indicates the expectation without this additional information—both are presented since this design capability of LSC is yet to be determined. Total statistics limit the outcome, and further running can proceed with IsoDAR and DUNE as it comes online if a NuTeV-like deviation is observed.
This result makes use of the $\bar{\nu}_{e}$ flux. An important design point is that the contemporaneous $\approx 1.7$M IBD events, which have a very well understood cross section, as discussed above, can be used to determine the flux normalization, allowing measurement of $\sin^{2}\theta_{W}$ with a precision of 1.9% (see Fig.7 left).
@@ 2.3.3 New Interactions
Figure 7: Left: Weak Mixing Angle expectation for IsoDAR compared to present results and the future DUNE expectation. Right: expectation for IsoDAR in NSI parameter space. From Ref. [10].
Example 2: Non-Standard Interactions of Neutrinos —
@@
The accurately predicted energy dependence of the ES sample enables a precise phenomenological search for non-standard neutrino interactions (NSIs). This powerful search for lepton-lepton couplings complements the lepton-quark coupling searches by the active, global coherent neutrino scattering program. The SM right- and left-handed couplings, $g_{R}$ and $g_{L}$, are altered by NSI corrections ($\epsilon_{ee}^{eR}$ and $\epsilon_{ee}^{eL}$) for electron flavor antineutrino interactions with electrons (hence “ee”) to $\tilde{g}_{R}=g_{R}+\epsilon_{ee}^{eR}$ and $\tilde{g}_{L}=g_{L}+\epsilon_{ee}^{eL}$. The SM expectation for IsoDAR@Yemilab in the parameter space of these deviations is seen in Fig. 7 (right), where red assumes background reduction using Cherenkov light, and green assumes only scintillation light reconstruction. This can be compared to the world result [28] in black, which is dominated by $\nu_{e}$–electron ES scattering from beam dump experiments.
This analysis uses the $\bar{\nu}_{e}$ flux. Because the IsoDAR result is very pure antineutrino, the expectation seen in Fig. 7 (right) is rotated with respect to the past results, where the fluxes had a mix of $\nu_{e}$ and $\bar{\nu}_{e}$ content. The result complements searches for NSIs from coherent scattering experiments, which aim to discover new couplings between neutrinos and nucleons. For a review of NSI searches, see Ref. [20].
3 Requirements for the Proton Driver
@@
As shown in Fig. 1, the Proton Driver consists of a 60 MeV/amu cyclotron accelerating and extracting 5 mA of $\mathrm{H}_{2}^{+}$, a transport line that strips the $\mathrm{H}_{2}^{+}$ ions close to the extraction point, analyzes the stripped beam to monitor the condition of the stripping foil, and transports the 10 mA of protons to the target. The target consists of three nested hemispherical beryllium shells, the outermost being 20 cm in diameter, cooled with heavy water flowing around the shell structure. The proton beam is shaped, with quadrupole magnets and wobbler magnets to produce a distribution that allows for optimal heat transfer to the cooling water. Neutrons produced in the target system are moderated and flood into a roughly-spherical sleeve, about 1 meter in diameter, containing a mixture of beryllium chips and highly-enriched 7Li. The beta-decay of the 8Li resulting from the capture of neutrons produces the neutrinos of interest in the experiment.
@@
The principal requirements for the accelerator system are to reliably produce a 10 mA, 60 MeV beam of protons in the most efficient manner possible with the lowest possible loss of particles in the acceleration process, particularly above the Coulomb threshold. The experience of the cyclotron system at PSI [29] is that keeping the total power of beam loss during cyclotron extraction to less than 200 W still allows access for component maintenance. Considering that the total beam power is 600 kW, this is a very stringent requirement. To be decisive within five years, we require a CW beam with 80% availability. Secondary requirements like compactness stem from the underground installation. Section 6 addresses the complex issue of transport and assembly of large and heavy cyclotron components in the underground environment at Yemilab.
@@ 3.1 Overview of the Accelerator Layout
Figure 8: Overview of the components of the HCHC-60 cyclotron system. Only one RF cavity is shown to not clutter the image. “Quads+SI” refers to focusing quadrupoles leading into the spiral inflector (SI). The beam is produced in the ion source at the top, shaped and guided into the RFQ by the LEBT, which is partially embedded in the cyclotron, bent onto the horizontal plane by the spiral inflector, and then acceleratred in the cyclotron.
To guide the reader through this PDR and, indeed, the path of an ion from genesis in the ion source to the full 60 MeV/amu and extraction, we provide an overview in Fig. 8. The main components of the system, which will be described in much detail in the following chapters, are:
  • The ion source, providing 10 mA of $\mathrm{H}_{2}^{+}$ at an emittance $0.1\pi-$$m$-mrad.
  • The Low Energy Beam Transport (LEBT), shaping and guiding the beam to the RFQ and providing diagnostics, and chopping for commissioning and personnel and machine protection.
@@
3.2 Status of the HCHC-60 Development
The HCHC-60 contains several subsystems that have never been used or combined in this way. To demonstrate the most critical aspects, we have begun an experimental campaign, building and testing some of these subsystems. Some already exist and are being tested, others are at the preliminary design level. Hence the overall status of the project is the PDR stage. The following gives a brief overview of the status of each subsystem.
  • Ion Source. We have designed and built two iterations of a multicusp ion source (MIST-1 and MIST-2). MIST-2 is currently being experimentally tested.
  • RFQ. The RFQ technical design is finished and it is currently being fabricated by Bevatech GmbH in Germany. We expect delivery in spring 2026 and will commission the RFQ together with the ion source.
  • HCHC-1.5 test cyclotron. A smaller version of the HCHC-60, this 1.5 MeV/amu test cyclotron is fully funded and currently in the Request For Proposal (RFP) stage. We expect vendor selection by the end of summer 2025.
  • HCHC-60. The full 60 MeV/amu cyclotron has reached the PDR stage. We describe the design and simulations in detail in the later sections of this document. We have not yet begun procurement of any hardware for the full cyclotron.
The ion source and RFQ could be reused directly in the final HCHC-60 cyclotron. The central region of the HCHC-1.5, while designed to mimic as much as possible the HCHC-60 central region, will have to be adapted to the magnetic field of the HCHC-60.
4 Design of the Front End
@@ 4.1 The Ion Source
The $\mathrm{H}_{2}^{+}$ beam to be injected into the cyclotron and accelerated to 60 MeV/amu will be produced by a filament-driven multicusp ion source. We have built a first prototype, the MIST-1 (Multicusp Ion Source at MIT - Version 1), the technical details and first measured results of which are reported in Refs. [30, 31, 32, 33, 34]. Here we summarize these results and present ongoing upgrade work (MIST-2). The upgrade to MIST-2 improves on MIST-1 by implementing refinements from lessons learned during the systematic measurements and runtime experiences described below. The design of the upgraded MIST-2 has been finalized and is in the final stages of fabrication and assembly. A publication on the MIST-2 design and performance is forthcoming.
@@ 4.1.1 Operating Principle
In a multicusp ion source [35], a plasma is confined by a set of permanent magnets that are arranged around the plasma chamber with alternating polarity, thereby creating “cusps” in the magnetic field. On average, the electrons and ions produced in the center of the chamber see an increasing magnetic field as they travel outward, which eventually turns them back. One notable exception are the cusps, which are the natural loss points of this configuration.
@@
[…]
Table 2: A summary of the parameters for the MIST-1 ion source.
@@ 4.1.2 Mechanical Design
Figure 9: Cross section of MIST-1 with extraction system, LEBT, and diagnostics. The diagnostics comprise an AC Current Transformer (ACCT) and a set of four capacitive button pickups right. Both are placed between the solenoid magnet and the RFQ at the end of the LEBT. Also shown is an example set of ion trajectories simulated with IBSimu and WARP. Adapted from Ref. [36].
For visual reference as we discuss the details of the ion source, 3D CAD renderings of MIST-1 (with subsequent LEBT) and MIST-2 are provided in Figs. 9 and 10, respectively. The main ion source parameters are also summarized in Table 2.
The MIST-type ion sources comprise a plasma chamber surrounded by a “water-jacket,” a back plate, and a front plate. The water jacket lets us pump de-ionized (DI) water around the plasma chamber for cooling. The radial confinement magnets are mounted inside the water-jacket and are cooled directly. At 6.5 cm, the plasma chamber is short compared to typical ion sources. This is a direct consequence of the short mean free path of $\mathrm{H}_{2}^{+}$ in the plasma/neutral gas of the ion source discussed above.
@@
The front plate holds the front confinement magnets as well as the plasma aperture, a W-Cu (75% - 25%) alloy (aka “Elkonite”) plate that is embedded in the front plate. We have several plasma apertures with different-size openings ranging from 4 mm to 8 mm to test beam quality and total extractable current for various aperture diameters.
As seen in Figs. 9 and 10, following the plasma aperture is the “extraction system” which shapes and guides the beam as it is initially formed. The MIST-1 extraction system comprises five electrodes which shape and accelerate the ion beam. These include a plasma electrode (+15 kV), a puller electrode (-3 kV to -20 kV), and three electrostatic lenses – Lens 1 (+10 kV to +15 kV), lens 2 (-3 kV to -15 kV), and lens 3 (+10 kV to +15 kV).
Figure 10: MIST-2 CAD rendering calling out the most important features.
MIST-2 versus MIST-1
@@
  • Replace all stainless steel components with oxygen-free high-conductivity (OFHC) copper to improve heat transfer
  • Increase the cross-sectional areas of all water cooling channels
@@
These changes will improve long-term running stability, reduce maintenance downtime, and allow for a better understanding of the plasma and beam conditions. Our high-fidelity simulations using IBSimu [37] and WARP [38] suggest that the beam dynamics downstream of the six-way cross will not change between the MIST-1 and MIST-2 configurations. All particle simulations and designs still hold for LEBT, RFQ, and cyclotron.